OpenStudy (anonymous):

PT tan (θ+30) + cot(θ-30) = 1/sin2θ - sin60.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Is that whole in denominator on right hand side ??

OpenStudy (anonymous):

dividing sign ?

OpenStudy (anonymous):

\[\tan(A +B) = \frac{\tan(A) + \tan(B)}{1 - \tan(A) \tan(B)}\]

OpenStudy (anonymous):

Your right hand side is this?? \[\frac{1}{\sin(2 \theta) - \sin(60)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\tan(\theta + 30) = \frac{\sqrt{3} \tan(\theta) + 1}{\sqrt{3} + \tan(\theta)}\]

OpenStudy (anonymous):

Because: \[\tan(30) = \frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Sorry: Negative sign will be there in denominator: \[\tan(\theta + 30) = \frac{\sqrt{3} \tan(\theta) + 1}{\sqrt{3} - \tan(\theta)}\]

OpenStudy (anonymous):

okay .. what about cot ?

OpenStudy (anonymous):

cot formula ?

OpenStudy (anonymous):

If you don't know the formula for cot then you can write that cot as 1/tan..

OpenStudy (anonymous):

Do you know the formula for cot ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay then can you use that ??

OpenStudy (anonymous):

cot (A+B) = cot A cot B - 1 / cot B+cot A correct ?

OpenStudy (anonymous):

Oh I think we are going in lengthy way..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

See I think if we will do this way then it may get easy..

OpenStudy (anonymous):

\[\tan(\theta) = \frac{\sin(\theta)}{\cos(\theta)}\]

OpenStudy (anonymous):

For simplicity now : Let theta + 30 = x and theta - 30 = y: \[\frac{\sin(x)}{\cos(x)} + \frac{\cos(y)}{\sin(y)} \implies \frac{\sin(x)\sin(y) + \cos(x) \cos(y)}{\cos(x) \sin(y)}\]

OpenStudy (anonymous):

\[\sin(x) \sin(y) + \cos(x)\cos(y) \implies \cos(x - y)\]

OpenStudy (anonymous):

y did yu delete ?

OpenStudy (anonymous):

\[x - y \implies \theta + 30 - \theta + 30 = 60\]

OpenStudy (anonymous):

I have used wrong formula for numerator.. now this one I have used is right. So the numerator is now : \[\cos(60) = \frac{1}{2}\]

OpenStudy (anonymous):

So the whole thing has now become: \[\implies \frac{1}{2\cos(x) \sin(y)}\] Getting till here??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Tell me the step you are not getting I will explain you then in far better way than this ??

OpenStudy (anonymous):

no i understand

OpenStudy (anonymous):

Are you sure you have got till here ??

OpenStudy (anonymous):

oh yes water ;)

OpenStudy (anonymous):

Can you solve the denominator now ??

OpenStudy (anonymous):

will try

OpenStudy (anonymous):

hold

OpenStudy (anonymous):

Ok, held on..

OpenStudy (anonymous):

am trying to do this sum frm d begining on my own by using the formula yu said , oly then will i first know if i have any doubt !

OpenStudy (anonymous):

Meaning ??

OpenStudy (anonymous):

You are going to solve it by using tan ????

OpenStudy (anonymous):

am just gonna try re-doin what you did just to make sure everything got in

OpenStudy (anonymous):

Okay.. But don't forget that you are going to solve for denominator part on your own.. For this part I mean: \[2 \cos(x) \sin(y)\]

OpenStudy (anonymous):

doubt !

OpenStudy (anonymous):

Where ??

OpenStudy (anonymous):

well this one , sinX sinY + cosX cosY = cos(X-Y)

OpenStudy (anonymous):

how cos(X-Y) ?

OpenStudy (anonymous):

I guessed the same.. ha ha ha..

OpenStudy (anonymous):

Do you know the formula for cos(x-y) ???

OpenStudy (anonymous):

\[\cos(x-y) = \cos(x) \cos(y) + \sin(x) \sin(y)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

See if x and y are greater than one another that will not change the formula because: \[\cos( - \theta) = \cos(\theta)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is your numerator now tell me ??

OpenStudy (anonymous):

what numerator ?

OpenStudy (anonymous):

The part where you had doubt write that part..

OpenStudy (anonymous):

sin(X) sin(Y) + cos(X) cos(Y) = cos(X-Y)

OpenStudy (anonymous):

\[\frac{\sin(x)}{\cos(x)} + \frac{\cos(y)}{\sin(y)} \implies \frac{\sin(x)\sin(y) + \cos(x) \cos(y)}{\cos(x) \sin(y)}\] Yeah here can you use that formula now ??

OpenStudy (anonymous):

water yu havn't yet solved my doubt .. i asked yu how cos(X-Y)

OpenStudy (anonymous):

@waterineyes ?

OpenStudy (anonymous):

It is a basic identity...

OpenStudy (anonymous):

i duno

OpenStudy (anonymous):

\[\cos(x + y) = \cos(x)\cos(y) - \sin(x) \sin(y)\] \[\cos(x-y) = \cos(x) \cos(y) + \sin(x) \sin(y)\]

OpenStudy (anonymous):

i kno this

OpenStudy (anonymous):

Can you remember these? I thought you know this..

OpenStudy (anonymous):

yea i kno

OpenStudy (anonymous):

You know this and you don't know that how? They are the same I think..

OpenStudy (anonymous):

how WATER ?

OpenStudy (anonymous):

Note that two identities in your notebook and remember those identities..

OpenStudy (anonymous):

Are you asking for the proof of cos(x-y) ??

OpenStudy (anonymous):

let me think .. what is happening

OpenStudy (anonymous):

These are basic identities you have to remember.. If you are asking for proof then I can tell you in detail next saturday I have to go now. For that last denominator part : use this formula: \[2\cos(x) \sin(y) = \sin(x + y) - \sin(x - y)\] After using this you will get right hand side..

OpenStudy (anonymous):

okay understood

OpenStudy (anonymous):

thanks @waterineyes

OpenStudy (anonymous):

Just try once you will get that we are close to right hand side... Bye take care ashna..

OpenStudy (anonymous):

Welcome..